If Sally can paint a house in 4 hours, and John can paint the same house in 6 hour, how long will it take for both of them to paint the house together?
Answer by ewatrrr(24785)


Hi
Let x represent the time it will take them together
Question States: Per hour being the Equalizer
1/4 + 1/6 = 1/x Mulitplying each term by 12x so as all denominators = 1
3x + 2x = 12
5x = 12
x = 12/5 = 2.4 hr or 2hr and 24min together.
CHECKING our Answer
3/12 + 2/12 = 5/12 = 1/(12/5)